### v1

# ConeGS: Error-Guided Densification Using Pixel Cones for Improved Reconstruction with Fewer Primitives

Abstract
3D Gaussian Splatting (3DGS) achieves state-of-the-art image quality and real-time performance in novel view synthesis but often suffers from a suboptimal spatial distribution of primitives. This issue stems from cloning-based densification, which propagates Gaussians along existing geometry, limiting exploration and requiring many primitives to adequately cover the scene. We present ConeGS, an image-space-informed densification framework that is independent of existing scene geometry state. ConeGS first creates a fast Instant Neural Graphics Primitives (iNGP) reconstruction as a geometric proxy to estimate per-pixel depth. During the subsequent 3DGS optimization, it identifies high-error pixels and inserts new Gaussians along the corresponding viewing cones at the predicted depth values, initializing their size according to the cone diameter. A pre-activation opacity penalty rapidly removes redundant Gaussians, while a primitive budgeting strategy controls the total number of primitives, either by a fixed budget or by adapting to scene complexity, ensuring high reconstruction quality. Experiments show that ConeGS consistently enhances reconstruction quality and rendering performance across Gaussian budgets, with especially strong gains under tight primitive constraints where efficient placement is crucial.
10k50k100k200k500k1M2M# Gaussians2424252526262727282829293030PSNR123OursMCMCEDGS
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Ours
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
MCMC [25]
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
EDGS [27]
Figure 1: ConeGS replaces cloning-based densification with a novel method that generates pixel-cone-sized primitives in regions of high image-space error. By improving placement and removing reliance on existing scene structure thanks to a flexible iNGP-based exploration, it achieves higher reconstruction quality than baselines using the same number of primitives. Results are averaged over Mip-NeRF 360 [3] and OMMO [33] datasets, with a visual comparison on the truck scene from Tanks & Temples [26].
22footnotetext: Corresponding author.
1 Introduction
Figure 2: Densification comparison. Cloning-based methods are difficult to tune, and the resulting primitives may require many iterations to fit correctly into the scene. ConeGS, by contrast, places primitives precisely using the pixel viewing cone size, enabling faster scene integration without reliance on the existing geometry.
Neural Radiance Fields (NeRF) [37] have significantly advanced novel view synthesis, achieving remarkable fidelity in scene reconstruction. However, representing scenes with neural networks makes NeRF slow to train and render, though it provides smooth parameterization and flexibility to handle changes in scene structure. Recently, 3D Gaussian Splatting (3DGS) [24] has gained attention as a faster, more practical alternative to NeRF, explicitly modeling scenes with sets of 3D Gaussians to achieve interactive rendering speeds while maintaining competitive visual fidelity. However, 3DGS increases expressiveness through cloning and splitting, which offer limited exploration, rely on hard-to-define densification rules, and generate many unnecessary primitives. As a result, primitives often accumulate in suboptimal regions, leaving large parts of the scene underrepresented or mispredicted.
To address these issues, we propose ConeGS, which replaces cloning-based densification with a novel strategy that targets pixels exhibiting high photometric error. By sampling these pixels and using depth estimates from a fast Instant Neural Graphics Primitives (iNGP) [39] reconstruction, new Gaussians are placed precisely in regions where the current representation is insufficient. This targeted placement increases expressiveness in areas requiring higher primitive density, improving reconstruction quality while avoiding redundant primitives. To determine the size of new Gaussians, we draw inspiration from Mip-NeRF [2]. During densification, each Gaussian is initialized according to the size of the viewing cone of the pixel from which it is generated at the specified depth. Their initial size is thus defined directly by their image-space coverage, eliminating the need for local size analysis or adjustments to reconstructed regions. Figure 2 illustrates the effectiveness of the proposed approach. Combined with a pre-activation opacity penalty that quickly removes redundant Gaussians, this enables scene representation with fewer primitives while preserving high reconstruction quality. We further incorporate two primitive budgeting strategies to regulate the total number of primitives, either through a fixed budget or by adapting to scene complexity. ConeGS outperforms baseline methods across diverse datasets and a wide range of primitive budgets. The advantage is most pronounced under tight primitive budgets. At higher budgets, it matches the quality of cloning-based methods, where efficient primitive placement is less critical, while still rendering faster than the baselines. In summary, our contributions are:
  • A densification strategy that places new Gaussians in regions of high photometric error in image space, guided by depth estimates from an iNGP-based geometric proxy.
  • An approach that determines the size of new Gaussians from the viewing cones of the pixels from which they are generated.
  • An improved opacity penalty that promptly removes low-opacity Gaussians, combined with a budgeting strategy that balances scene complexity and primitive count.
Finally, our method is also compatible with other 3DGS improvements, making it straightforward to integrate with existing approaches for greater efficiency, or with methods where cloning strategies are ambiguous or hard to formalize [34, 20, 49].
2 Related work
Neural Radiance Fields: NeRFs [37] represent scenes as continuous volumetric radiance fields, enabling high-quality novel view synthesis. This is achieved by parameterizing the scene with a neural network (typically an MLP), whose weights encode the scene globally. Despite producing photorealistic results, these methods rely on costly volumetric rendering and remain computationally inefficient. Extensions such as Mip-NeRF [2] and Mip-NeRF360 [3] reduce aliasing via conical frustum integration, while Zip-NeRF [4] improves view consistency with hierarchical sampling and multi-scale supervision. Hybrid approaches [45, 56, 46, 7] mitigate this by combining explicit data structures with compact neural representations, enabling faster optimization and real-time rendering. Instant Neural Graphics Primitives (iNGP) [39] further accelerate training through multi-resolution hash-grid encoding and shallow MLPs.
Primitive-based Differentiable Rendering: 3D Gaussian Splatting (3DGS) [24] has emerged as an efficient alternative to Neural Radiance Fields (NeRF) [37]. Rather than modeling the scene as a global volume, 3DGS represents it with local explicit 3D Gaussians and uses differentiable rasterization, resulting in significantly faster rendering. Its balance of fidelity and efficiency has attracted significant attention and spurred a wide range of follow-up research. Prior works have focused on tackling anti-aliasing [57, 53], reconstructing dynamic scenes [52, 54], enabling generative content creation [48, 66], reducing rendering artifacts [43], substituting alpha composition with volumetric rendering [35, 47], extracting geometry [21, 58, 16], level-of-detail reconstruction [44], frequency-based regularization [60, 59], and introducing new primitives or kernel functions [18, 20, 31, 49]. Recent efforts have also targeted reducing computational and memory costs, often through feature quantization or code-book encoding [41, 15, 10, 34], or scene simplification [63]. [11] reduces computation by lowering the number of primitives through an aggressive densification and pruning strategy, while [12, 11, 64] insert new Gaussians at the currently estimated depths using re-initialization. Unlike our method, this approach overwrites existing structures instead of adding new points, and further depends on the scene already being well reconstructed. [25] improves primitive distribution and exploration by incorporating positional errors and applying penalties to opacity and scaling. Closely related to our approach, several works focus on improving densification to reduce redundancy and better capture fine details. Strategies include refining cloning heuristics [25, 5, 22], per-Gaussian property- or saliency-based cues [36], geometry- and volume-aware criteria [23, 1, 65], addressing gradient collision [55], perceptual sensitivity [64], learnable schemes [40, 32], and based on Gaussian Processes [17]. Some works target densification in challenging settings [38], filling holes in the representation [9, 29], though typically adding only a few primitives. PixelSplat [6] models dense probability distributions for more robust Gaussian placement, influencing later approaches [8]. Recent work [27] suggests that densification may be unnecessary for high-quality reconstruction given strong initialization. Like our method, they start by estimating scene geometry, but rely only on correspondences from a pretrained dense matching network, without enhancing densification, and at higher GPU memory cost than our approach. Concurrent work [62], employs Gaussians with spatially varying texture colors, improving fine-detail reconstruction and reducing the number of primitives needed. Other methods use neural radiance fields for depth supervision [14, 28] or point cloud extraction [42, 14, 50] to initialize a scene, but not to improve densification directly. Concurrent work [13] applies NeRF for initialization and limited densification, constrained by existing Gaussian locations, and does not explore varying Gaussian sizes, which we find beneficial for reconstruction quality.
3 Preliminaries
3D Gaussian Splatting: 3DGS [24] represents a scene as an unordered set of 3D Gaussian primitives $\{\mathcal{G}_{i}|i=1,\ldots,M\}$. Each primitive $\mathcal{G}_{i}=(\mathbf{p}_{i},\mathbf{s}_{i},\mathbf{R}_{i},o_{i},\mathbf{c}_{i})$ is defined by its position $\mathbf{p}_{i}\in\mathbb{R}^{3}$, scaling vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$, rotation matrix $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$, opacity $o_{i}\in\mathbb{R}$, and view-dependent color $\mathbf{c}_{i}\in\mathbb{R}^{3}$. The color $\mathbf{c}_{i}$ is represented by spherical harmonics (SH) coefficients $\mathbf{k}_{i}\in\mathbb{R}^{3L}$, where $L$ is the number of coefficients determined by the chosen SH order. The 3D covariance matrix is given by $\mathbf{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{T}\mathbf{R}_{i}^{T}$, where $\mathbf{S}_{i}=\text{diag}(\mathbf{s}_{i})$ is the scaling matrix. The color $\hat{C}$ of a pixel is computed by $\alpha$-blending over a set of $N$ Gaussians, sorted by depth, whose projections overlap the pixel:
$\hat{C}$ $=\sum\nolimits_{i\in N}\mathbf{c}_{i}\alpha_{i}\prod\nolimits_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
$\alpha_{i}$ $=o_{i}K(\mathbf{p}_{c},\boldsymbol{\mu}^{\text{2D}}_{i},\mathbf{\Sigma}^{\text{2D}}_{i}),$ (2)
where $\alpha_{i}$ is the blending weight of the $i$-th Gaussian, $\mathbf{p}_{c}$ is the pixel center in image coordinates, $\boldsymbol{\mu}_{i}^{\text{2D}}$ and $\mathbf{\Sigma}_{i}^{\text{2D}}$ are the 2D projected mean and covariance of $\mathcal{G}_{i}$, and $K(\cdot)$ is a Gaussian filter response in screen space. The exact form of $K$ depends on the chosen filter [24, 67, 57]. Gaussians are traditionally initialized from an SfM point cloud, with each component of $\mathbf{s}_{i}$ set equal to the mean Euclidean distance to the three nearest neighbors $\mathcal{N}_{3}(i)$ of Gaussian $i$:
$\mathbf{s}_{i}=(s_{i},s_{i},s_{i}),\quad s_{i}=\frac{1}{3}\sum_{k\in\mathcal{N}_{3}(i)}\|\mathbf{p}_{k}-\mathbf{p}_{i}\|\quad.$ (3)
During training, the Gaussian parameters are optimized with the loss:
$\mathcal{L}_{\text{GS}}=(1-\lambda)\,\text{MAE}(I,I^{*})+\lambda\,\mathcal{L}_{D\text{-}SSIM},$ (4)
where $\lambda=0.2$, MAE is the mean absolute error between the rendered image $I$ and the ground-truth image $I^{*}$, and $\mathcal{L}_{D\text{-}SSIM}=1-\text{SSIM}(I,I^{*})$ [51].
Neural Radiance Fields: NeRFs [37] model a scene as a continuous 3D field that maps a 3D location along a camera ray and the viewing direction of the corresponding pixel to a density $\sigma\in\mathbb{R}$ and color $\mathbf{c}\in\mathbb{R}^{3}$. A camera ray is parameterized as $\mathbf{r}(t)=\mathbf{p}_{\text{cam}}+t\mathbf{d}$, where $\mathbf{p}_{\text{cam}}$ is the camera position and $\mathbf{d}$ is a unit direction vector pointing toward the center of a pixel. Each ray is discretized into $N$ intervals defined by distances $\{t_{i},t_{i+1}\}_{i=1}^{N}$. For each sample position $\mathbf{r}(t_{i})$ along the ray, the NeRF is queried to predict the sample’s color $\mathbf{c}_{i}$ and density $\sigma_{i}$. Using volumetric rendering [37], the corresponding pixel color is approximated as:
$\hat{C}$ $=\sum\nolimits_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod\nolimits_{j=1}^{i-1}(1-\alpha_{j}),$ (5)
$\alpha_{i}$ $=1-\exp(-\sigma_{i}\delta_{i})\quad\text{with}\kern 5.0pt\delta_{i}=t_{i+1}-t_{i}.$ (6)
Here, $\alpha_{i}$ is the opacity of the $i$-th sample, $\delta_{i}$ is the length of its ray segment, and the product term represents the transmittance $\tau_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$.
Sampling only a single ray per pixel can lead to blur and aliasing. Mip-NeRF [2] addresses this by replacing the ray with a cone that models the pixel footprint, i.e. the 3D volume a pixel covers in world space. The cone is divided into frustums, and integration is performed over these volumes rather than along a 1D line. The cone’s radius $r_{\mathrm{cone}}(t)$ defines the cross-section of the pixel cone at distance $t$ and is computed from the directions of rays passing through the pixel and its neighbors:
$r_{\mathrm{cone}}(t)=t\;\frac{\|\mathbf{d}_{x}-\mathbf{d}\|+\|\mathbf{d}_{y}-\mathbf{d}\|}{2},$ (7)
where $\mathbf{d}$ is the direction of the ray through the center of the pixel, and $\mathbf{d}_{x},\mathbf{d}_{y}$ are the directions of rays through the neighboring pixels in the $x$ and $y$ directions, respectively.
(a) Initialization (Section 4.1)
(b) Optimization (Section 4.2)
Figure 3: Overview of the ConeGS pipeline. ($a$) First, an iNGP reconstruction is obtained to serve as a geometric proxy for object surfaces, guiding the placement of Gaussians both during scene initialization and throughout the 3DGS optimization process. ($b$) During 3DGS optimization, ConeGS performs error-guided densification by sampling a subset of pixels with high $L_{1}$ error. For each sampled pixel, a new Gaussian $\mathcal{G}$ is created along the pixel’s viewing cone at the depth estimated by iNGP and scaled to match the cone’s size. New Gaussians are accumulated and, every 100 iterations, inserted into the scene after pruning those with low opacity. Blue arrows indicate gradient updates to Gaussian parameters, and the red arrow marks scene updates.
4 Method
This section outlines our ray-based densification approach for 3DGS. First, we explain how we use an iNGP model as a geometric proxy to initialize the 3D Gaussian scene (Section 4.1). Next, we detail our ray-based densification strategy, which uses the iNGP to place pixel-cone-sized Gaussians in high-error regions, along with associated optimization changes (Section 4.2). Finally, we provide additional implementation details in Section 4.3. An overview of the complete pipeline is shown in Figure 3.
4.1 Initialization
We use a trained iNGP model [39] as a geometric proxy to initialize the 3DGS scene and guide densification. Trained briefly on input images, it provides accurate depth estimates, that position both the initial Gaussian primitives and those added later during densification, with minimal impact on training time. Additionally, the depths can be evaluated on the fly during optimization, reducing both memory usage and computation compared to precomputing all depth maps. We initialize the scene with $\mathcal{P}_{\text{init}}$ Gaussians, set to one million as in [42], or fewer if a smaller budget is specified (see Section 4). To construct this set, we uniformly sample $\mathcal{P}_{\text{init}}$ image-pixel pairs $(I,u,v)$ from the training set pixel domain $\mathcal{I}_{\text{train}}$. Each sampled image-pixel pair defines exactly one Gaussian in the initialized scene. For each sample, we define its associated camera ray
$\mathbf{r}_{I}(u,v,t)=\mathbf{p}_{I}+t\,\mathbf{d}_{I}(u,v),$ (8)
where $I$ is an image index, $(u,v)$ are pixel coordinates, $\mathbf{p}_{I}\in\mathbb{R}^{3}$ is the camera center, and $\mathbf{d}_{I}(u,v)\in\mathbb{R}^{3}$ is the normalized ray direction. We query the iNGP along this ray to obtain discrete transmittance values $\{\tau_{k}\}$, from which the median depth $t_{\text{med}}$ is computed as:
$t_{\text{med}}=t_{k}\quad\text{where}\quad\tau_{k-1}>0.5\geq\tau_{k}.$ (9)
The center of the $j$-th Gaussian primitive is then set to:
$\mathbf{p}_{j}=\mathbf{p}_{I}+t_{\text{med}}\,\mathbf{d}_{I}(u,v),$ (10)
yielding the set of initial centers $\{\mathbf{p}_{j}\}_{j\in\mathcal{I}{\text{sample}}}$ with $\mathcal{I}_{\text{sample}}\subset\mathcal{I}_{\text{train}}$. The scale $\mathbf{s}_{j}$ is initialized isotropically using the average distance to the three nearest neighbors $\mathcal{N}_{3}(j)$, following Eq. (3). The rotation is set to identity $\mathbf{R}_{j}=\mathbf{I}$, the opacity to $o_{j}=0.1$, and the SH coefficients to:
$\mathbf{k}_{j}=\left(\mathbf{k}_{1:3,j},\mathbf{k}_{4:L,j}\right),\quad\mathbf{k}_{1:3,j}=\mathbf{c}_{j,0},\quad\mathbf{k}_{4:L,j}=\mathbf{0},$ (11)
where $\mathbf{c}_{j,0}$ is the RGB color for the sampled pixel $(I,u,v)$ rendered with iNGP using a zeroed-out view direction. Although our densification strategy can achieve high-quality reconstructions without scene initialization, we retain this step to ensure consistently strong performance across all metrics (see Section 5.2).
4.2 Optimization
We fully replace the standard 3DGS cloning-based densification with an error-guided strategy that adapts the iNGP-based ray-depth rendering procedure from Section 4.1 to position new Gaussian primitives. Below, we outline the sampling, scaling, budgeting, and pruning stages of our densification pipeline.
Error-Weighted Gaussian Densification: To limit the number of primitives while targeting poorly reconstructed regions, we add new Gaussians at pixels with high photometric error. At iteration $j$, we render an image $I_{j}$ and compute the per-pixel absolute error ($L_{1}$ loss) $E(\mathbf{p})=|I_{j}(\mathbf{p})-I^{*}(\mathbf{p})|$ with respect to the ground-truth image $I^{*}$. We then sample $N_{\text{sample}}$ pixels without replacement according to a multinomial distribution $\mathcal{M}$ with probabilities proportional to the normalized error map:
$\{\mathbf{p}_{s}\}_{s=1}^{N_{\text{sample}}}\sim\mathcal{M}\left(N_{\text{sample}},\frac{E(\mathbf{p})}{\sum_{\mathbf{p}^{\prime}\in I_{j}}E(\mathbf{p}^{\prime})}\right),$ (12)
While $L_{1}$ loss does not always indicate possible improvements and can also arise from noise or difficult-to-optimize reflections, we found it to be a reliable indication of lacking expressiveness, especially at low primitive budgets. For each sampled pixel $\mathbf{p}_{s}$, a new Gaussian $\mathcal{G}_{s}$ is created, with its center placed along the corresponding ray at the median depth $t_{\text{med},s}$ given by the iNGP. Newly spawned Gaussians are appended to an accumulation set:
$\mathcal{G}_{\text{accum}}\leftarrow\mathcal{G}_{\text{accum}}\cup\left\{\mathcal{G}_{s}\right\}_{s=1}^{N_{\text{sample}}}.$ (13)
Every 100 iterations, low-opacity Gaussians in the scene $\mathcal{G}_{\text{scene}}$ are pruned, and Gaussians in $\mathcal{G}_{\text{accum}}$ are merged into the scene:
$\mathcal{G}_{\text{scene}}\leftarrow\mathcal{G}_{\text{scene}}\cup\mathcal{G}_{\text{accum}},\quad\mathcal{G}_{\text{accum}}\leftarrow\emptyset.$ (14)
The newly inserted Gaussians are then jointly optimized with existing primitives. Unlike cloning-based methods, which constrain new primitives to the vicinity of existing ones and thus hinder exploration of unseen regions, our approach places Gaussians directly in high-error areas, enabling effective scene coverage even far from existing geometry. Moreover, we preserve the integrity of well-reconstructed areas, since new Gaussians are added on top of the existing structure rather than created through splitting or cloning. Although occasional iNGP depth inaccuracies may introduce misplaced Gaussians, diverse viewpoint coverage ensures that inconsistent ones are quickly corrected or pruned, while multiview-consistent ones are retained. The full densification pipeline is shown in Figure 4.
Pixel-Footprint-Aligned Scaling: Selecting an appropriate scale for newly added Gaussians during densification is crucial. If primitives are too large, they may obscure fine details and be pruned prematurely, whereas overly small ones contribute little to the rendered image, yielding weak gradients and slowing convergence. Although $k$-NN–based scaling is effective for initialization, recomputing nearest-neighbor distances at every densification step is computationally expensive and sensitive to outliers. Large distances can produce inflated scales, causing new Gaussians to overlap well-reconstructed regions and hinder further optimization (see Section 5.2). To avoid these issues, we set the initial scale of each newly added Gaussian directly from the pixel footprint at the median depth $t_{\text{med},i}$ along its corresponding camera ray, as defined in Eq. 7:
$\mathbf{s}_{i}=\lambda_{\text{scale}}\,r_{\mathrm{cone}}(t_{\mathrm{med},i})\,(1,1,1),$ (15)
where $\lambda_{\text{scale}}=2$ converts the cone radius $r_{\mathrm{cone}}(t_{\mathrm{med},i})$ to the diameter of its cross-section. This ensures that, from the spawning viewpoint, the Gaussian’s projection onto the image plane approximately matches the pixel width, independent of scene depth. The assigned scale is only an initial value, with subsequent optimization steps jointly updating all primitives to allow newly added Gaussians to adjust to the existing scene. Our pixel-aligned, depth-aware scaling provides three key benefits: (1) it is independent of the current primitive distribution, avoiding the structural biases of cloning-based methods that replicate and reinforce local geometry, (2) pixel size allows Gaussians to contribute to optimization immediately and efficiently fit fine details while minimizing overlap with existing structure, and (3) its isotropic shape promotes stable multi-view integration, without shapes that the cloned Gaussians inherit.
Figure 4: Densification overview. Illustration of the proposed error-guided strategy. We render an image with 3DGS, compute the per-pixel $L_{1}$ error, sample pixels proportionally to their error magnitude, and place new Gaussians at the iNGP-predicted depth along the corresponding viewing rays.
Primitive Budgeting: We consider two budgeting strategies for controlling the number of Gaussians in the scene. The first enforces a hard upper bound, as in [25], ensuring that densification never exceeds the prescribed budget. This constraint regulates memory and computation while preventing uncontrolled growth of the primitive set. At each densification step, we set the number of sampled pixels $N_{\text{sample}}$ so that newly added Gaussians replace those pruned, avoiding excess primitives that would otherwise be discarded under the budget. This is computed as:
$N_{\text{sample}}=\frac{\max(0.2N_{\text{GS}},\ 1.2N_{\text{last}})}{100},$ (16)
where $N_{\text{GS}}$ is the current total number of Gaussians, and $N_{\text{last}}$ is the number of primitives inserted in the previous densification step, and the division by 100 reflects the densification interval. This formulation keeps $N_{\text{GS}}$ close to the budget limit even under aggressive pruning, maintaining consistent scene coverage throughout optimization.
The second strategy adapts the number of primitives to the scene’s complexity, enabling controlled growth without imposing a fixed upper bound:
$N_{\text{sample}}=\frac{\beta N_{\text{GS}}}{100}.$ (17)
Here, $\beta$ controls the growth rate of the primitive set. Smaller values balance the number of Gaussians added with those pruned, maintaining a relatively stable primitive count, whereas larger values yield higher primitive counts, increasing geometric detail at the cost of memory and computing power. With scene initialization at 1M primitives and the application of the opacity penalty, we balance Gaussians added and pruned, unlike [5], which requires a predefined upper limit on primitives.
Opacity-Regularized Pruning: Following [24, 25], Gaussians with opacity below $0.005$ are pruned every 100 iterations to remove primitives with negligible contribution to the rendered image. Earlier work has promoted sparsity through different strategies: periodically resetting opacities [24], which can destabilize training [5], reducing opacities by a constant amount after each densification [5], or introducing a post-activation opacity penalty $\mathcal{L}_{\mathbf{o}}^{\text{post}}=\|\sigma(\mathbf{o}_{\text{pre}})\|_{1}$ [25], where $\mathbf{o}_{\text{pre}}$ denotes the opacity logits before the sigmoid and $\sigma$ is the sigmoid function. This applies the strongest constraint around $0.5$ and only a weak penalty near the pruning threshold. In contrast, we employ a pre-activation opacity penalty $\mathcal{L}_{\mathbf{o}}^{\text{pre}}=\|\mathbf{o}_{\text{pre}}\|_{1}$. It provides a steady constraint across the full opacity range, including very low values, gradually reducing under-contributing primitives. The penalty acts throughout training, and our densification strategy can freely add new primitives, allowing any structure lost through pruning to be recovered more easily than with cloning-based approaches. In all experiments, this loss is scaled by $\lambda_{\mathbf{o}}=0.0002$.
4.3 Implementation Details
For the iNGP model, we use the proposal-based implementation from NerfAcc [30], trained for 20k iterations with the original setup and architecture. Gaussian optimization runs for 30k iterations, with our densification active for the first 25k. Unlike 3DGS, all SH components are optimized from the start, enabled by the stable initialization that removes the need for gradual SH introduction.
Mip-NeRF360 [3] OMMO [33] Tanks & Temples [26] DeepBlending [19]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Number of Gaussians limited to 100100k
3DGS [24] (SfM init.) 23.61 0.693 0.413 26.45 0.820 0.296 22.38 0.774 0.333 24.65 0.827 0.412
Foroutan et al. [14] 26.64 0.781 0.318 26.89 0.829 0.276 22.48 0.766 0.341 25.31 0.822 0.418
MCMC [25] (rand. init.) 25.72 0.730 0.369 25.92 0.808 0.313 21.45 0.750 0.365 27.94 0.859 0.369
MCMC [25] (SfM init.) 27.06 0.800 0.303 27.01 0.841 0.266 22.50 0.780 0.332 28.94 0.876 0.333
MCMC [25] (iNGP init.) 27.35 0.797 0.299 26.95 0.837 0.265 22.69 0.775 0.326 29.02 0.872 0.337
GaussianPro [9] 25.57 0.766 0.338 26.14 0.822 0.289 20.59 0.757 0.348 28.15 0.870 0.342
Perceptual-GS [64] 25.66 0.774 0.320 26.13 0.819 0.292 20.76 0.759 0.347 28.32 0.874 0.338
EDGS [27] 27.09 0.798 0.296 26.99 0.838 0.261 22.32 0.777 0.324 28.43 0.872 0.337
Ours 27.74 0.809 0.285 27.59 0.852 0.243 23.12 0.791 0.310 29.44 0.880 0.328
Number of Gaussians limited to 500500k
3DGS [24] (SfM init.) 28.22 0.821 0.260 28.96 0.883 0.196 23.54 0.816 0.265 29.25 0.882 0.302
Foroutan et al. [14] 28.88 0.862 0.204 28.80 0.884 0.185 23.52 0.816 0.257 29.61 0.886 0.297
MCMC [25] (rand. init.) 28.38 0.844 0.237 28.23 0.874 0.212 22.96 0.808 0.279 28.84 0.875 0.315
MCMC [25] (SfM init.) 28.82 0.861 0.214 28.72 0.885 0.194 23.68 0.825 0.259 29.44 0.887 0.308
MCMC [25] (iNGP init.) 29.02 0.867 0.198 28.75 0.885 0.186 23.57 0.823 0.243 29.61 0.885 0.288
GaussianPro [9] 28.02 0.827 0.253 28.32 0.876 0.206 22.31 0.802 0.286 29.33 0.886 0.301
Perceptual-GS [64] 28.66 0.856 0.211 28.42 0.876 0.203 22.77 0.813 0.267 29.44 0.888 0.296
EDGS [27] 28.82 0.865 0.193 29.01 0.893 0.168 23.47 0.831 0.229 29.33 0.889 0.286
Ours 29.08 0.870 0.190 29.14 0.892 0.170 23.69 0.829 0.229 29.86 0.891 0.285
Number of Gaussians bounded by Mini-Splatting2 [11]
Mini-Splatting2 [11] 28.89 0.875 0.183 28.06 0.875 0.198 22.79 0.823 0.239 29.99 0.898 0.279
Ours 29.26 0.875 0.179 28.90 0.887 0.179 23.66 0.829 0.231 29.82 0.891 0.280
  • $\dagger$
    Original code was not publicly available. Our implementation uses iNGP initialization and does not include the additional depth-based loss.
Table 1: Quantitative results with a Gaussian number limit of 100k and 500k. Mini-Splatting2 [11] does not support constraining the number of Gaussians during reconstruction, so we match its Gaussian count. We highlight the best, second best and third best results among methods with the same Gaussian counts. Per-scene metrics for selected methods are provided in the supplementary material.
5 Evaluation
EDGS [27] MCMC [25] Ours GT
Room [3] (50k)
Refer to caption Refer to caption Refer to caption Refer to caption
Dr Johnson [19] (100k)
Refer to caption Refer to caption Refer to caption Refer to caption
05 [33] (200k)
Refer to caption Refer to caption Refer to caption Refer to caption
13 [33] (500k)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative results comparing our method with MCMC [25] (with SfM point cloud initialization) and EDGS [27] on the Mip-NeRF 360 [3], OMMO [33], and DeepBlending [19] datasets, with varying Gaussian budgets (given in parentheses).
10k50k100k200k500k1M2M22.524.025.527.028.530.0PSNRBudget441k548k750k1.27M2.19M22.524.025.527.028.530.0No budget10k50k100k200k500k1M2M0.160.240.320.400.48LPIPS441k548k750k1.27M2.19M0.160.240.320.400.4810k50k100k200k500k1M2M# Gaussians100200300400500FPS441k548k750k1.27M2.19M# Gaussians100200300400500ConeGSMCMC (SfM init.)$MC$ (rand. init.)EDGS
Figure 6: PSNR, LPIPS and FPS plots for a specified primitive budget (left) and without a budget (right), where the number of primitives corresponds to the chosen $\beta$ values. Numerical results are provided in the appendix. Values are averaged across Mip-NeRF360 [3] and OMMO [33].
Dataset and Metrics: We evaluate our method on publicly available scenes from Mip-NeRF360 [3] and OMMO dataset [33], with 01 scene from OMMO resized to have 1600 pixels width. Following [24, 25], we also include the train and truck scene from Tanks & Temples [26], as well as Dr Johnson and playroom from the DeepBlending [19] dataset. We report PSNR, SSIM [51], and LPIPS [61], with rendering speeds averaged across the full test set. All FPS measurements were recorded on an NVIDIA RTX 2080 Ti, whereas training speeds are reported on an NVIDIA A100, since EDGS requires more memory. These training speeds do not include later-added speed improvements [36].
Baselines: We primarily compare our method against 3DGS [24], it’s extension with iNGP point cloud initialization [14], MCMC [25] using different initialization types (random, SfM, iNGP point clouds), as well as, GaussianPro [9], Perceptual-GS [64], EDGS (with densification) [27], with the densification stopped for all of them if the primitive budget is reached. If the number of primitives at initialization would be higher than the specified budget, the number of primitives is sampled uniformly to fit below it. In the random initialization settings we follow the process described in MCMC [25]. We additionally test on Mini-Splatting2 [11] by matching their final number of primitives instead of a specific budget, due to their method relying on generating a high number of initial Gaussians.
Ours
10k iters
Ours
20k iters
Ours
40k iters
EDGS [27]
3DGS [24]
(SfM init.)
MCMC [25]
(rand. init.)
MCMC [25]
(SfM init.)
PSNR $\uparrow$ 29.33 29.37 29.37 29.18 28.71 28.98 29.23
SSIM $\uparrow$ 0.877 0.88 0.881 0.877 0.846 0.865 0.875
LPIPS $\downarrow$ 0.171 0.168 0.166 0.168 0.222 0.204 0.190
FPS $\uparrow$ 134 137 137 131 112 92 95
3DGS time $\downarrow$ 20.7 20.5 20.5 19.4 22.6 26.3 25.1
Init. time $\downarrow$ 1.6 3.1 6.1 2.3 - - -
Overall time $\downarrow$ 22.3 23.6 26.6 21.7 22.6 26.3 25.1
Table 2: Quantitative results showing overall scene reconstruction timings for 3DGS [24], MCMC [25] and EDGS [27] methods, as well as our method with differing iNGP reconstruction durations, with the number of Gaussians capped at 11M. Results are averaged over the Mip-NeRF360 [3] dataset. 3DGS time reports the target scene optimization, including densification, while init. time shows the iNGP reconstruction needed by our method and the initial matching for EDGS [27].
5.1 Results
We observe improvements over the baselines across a wide range of specified budgets in Table 1, with plots comparing the most important methods on the budget and no-budget scenario in Figure 6. For a lower limit on Gaussians, we outperform the benchmarks across all datasets and metrics, while providing a competitive reconstruction quality compared to the best performing baselines on the high budget scenarios. In Table 2 we additionally show that even on a high number of primitives of 1M and including the iNGP training, our method provides competitive speed to other methods. The qualitative results in Figure 5 show significant improvement using a wide range of primitive budgets, demonstrating that for the same limit of primitives our method is able to produce a much better reconstruction, especially in areas that are challenging to properly capture on a low budget, such as isolated or high frequency structures. We provide additional qualitative and quantitative results, along with further scene analysis, in the appendix.
5.2 Ablations
We analyze the impact of individual components in our method by conducting a series of ablation studies, presented in Table 3. (a), (b) Longer iNGP reconstruction only slightly improves reconstruction quality. (c) continuing training the iNGP model also during optimization, (d) initializing Gaussians with the ground truth pixel color, or (e) predicting spherical harmonics with iNGP, leads to marginally worse results. Demonstrating the strength of our densification method, using (f) pixel-cone-sized primitives during initialization, or even not using any initialization (g), results in worse PSNR but maintains low LPIPS and considerably improves rendering speed, thanks to less overlap between primitives, reducing blending. (h) Sampling pixels uniformly instead of guiding Gaussian creation using the $L_{1}$ loss from the training set produces lower reconstruction quality, although due to uniform sampling in image space still focusing more on parts of the scene seen most across views, the drop is not drastic. Densifying with 3DGS depth (i), also without using iNGP even for initialization (j), strongly affects the results. Similarly, k-NN sizing of newly added primitives based on their closest neighbors (k), or changing the opacity penalty (l), (m), (n), has a large effect on reconstruction quality, reinforcing the benefits of our densification approach. (o), (p) Altering the size of Gaussians created during densification from their default pixel-width cone size slightly reduces reconstruction quality, although the small difference suggests that Gaussians are quickly resized to fit the scene.
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
Ours 27.74 0.810 0.285 328
(a) 10k iNGP iter. 27.74 0.808 0.287 313
(b) 40k iNGP iter. 27.72 0.811 0.284 333
(c) Train iNGP during 3DGS 27.73 0.809 0.285 310
(d) Color from GT image 27.73 0.807 0.287 320
(e) Prediction of SH with iNGP 27.74 0.805 0.290 320
(f) Cone-sized initialization 27.38 0.806 0.287 437
(g) Without initialization 27.46 0.811 0.285 415
(h) Uniform image-space sampling 27.51 0.806 0.286 307
(i) Densify with 3DGS depth 27.43 0.797 0.296 332
(j) SfM initialization + 3DGS depth dens. 27.15 0.790 0.302 329
(k) Densify with k-NN scaling 27.54 0.802 0.295 299
(l) No opacity penalty 27.31 0.794 0.301 294
(m) Post-densification opacity decrease [5] 27.46 0.798 0.297 256
(n) MCMC-style opacity penalty [25] 27.49 0.803 0.293 239
(o) $\lambda_{\text{scale}}=1$ 27.70 0.810 0.285 329
(p) $\lambda_{\text{scale}}=4$ 27.63 0.808 0.285 329
Table 3: Ablation study of our method with 100100k Gaussians, averaged over the Mip-NeRF360 [3] dataset. We highlight the best, second best, and third best results.
6 Conclusion
We introduce ConeGS, a reconstruction pipeline that replaces cloning-based densification with a method guided by photometric error and a coarse iNGP proxy, where new primitives are sized by pixel cones. Together with an improved opacity penalty, this allows creating primitives independently of existing structures through more flexible exploration. ConeGS consistently improves reconstruction quality and rendering performance across Gaussian budgets, with strong gains under tight primitive constraints. It achieves up to a 0.6 PSNR increase and a $20\%$ speedup over cloning-based baselines.
Limitations: ConeGS works well on standard scenes but can struggle with very large scenes, inaccurate camera poses, or sparse viewpoints, sometimes producing extra floating Gaussians (see appendix). Its gains are also limited at very high Gaussian budgets, where dense coverage reduces the benefit of error-guided placement, offering mainly faster rendering.
Appendix
This appendix introduces additional results (Section A1) and ablations (Section A2). We then discuss our reconstructed scene structure (Section A3), possible failure cases (Section A4), and provide visualizations for the different types of initializations mentioned in the main paper (Section A5). Finally, we discuss the parallels between 3DGS and NeRF rendering, which allows training a radiance field using the rendering from 3DGS (Section A6).
Appendix A1 Detailed results
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ # Gaussians
Ours ($\beta=0$) 29.12 0.873 0.183 185 542k
MCMC (SfM) 28.86 0.865 0.209 136
EDGS 28.87 0.868 0.187 194
Ours ($\beta=0.01$) 29.25 0.877 0.174 169 674k
MCMC (SfM) 29.06 0.870 0.199 119
EDGS 28.97 0.873 0.178 167
Ours ($\beta=0.02$) 29.34 0.880 0.166 146 942k
MCMC (SfM) 29.15 0.876 0.189 101
EDGS 29.10 0.878 0.167 134
Ours ($\beta=0.04$) 29.38 0.881 0.161 105 1.66M
MCMC (SfM) 29.32 0.882 0.174 77
EDGS 29.19 0.881 0.159 102
Ours 29.35 0.879 0.159 80 2.57M
MCMC (SfM) 29.56 0.887 0.164 55
EDGS 29.37 0.884 0.153 67
3DGS 29.03 0.870 0.184 69
Table A1: No-budget Mip-NeRF360. Comparison of reconstruction quality without a fixed limit on the number of primitives. For each cell, our method is run with a different $\beta$ parameter, which determines the number of Gaussians generated, and the other methods are limited to this number. In the last cell, the number of Gaussians is set to match the amount produced by 3DGS, with all other methods constrained accordingly. Results are averaged over the Mip-NeRF360 [3] dataset.
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ # Gaussians
Ours ($\beta=0$) 28.54 0.879 0.195 253 352k
MCMC (SfM) 28.37 0.877 0.207 199
EDGS 28.62 0.885 0.183 219
Ours ($\beta=0.01$) 29.02 0.890 0.176 237 438k
MCMC (SfM) 28.55 0.882 0.199 175
EDGS 28.84 0.890 0.173 211
Ours ($\beta=0.02$) 29.19 0.894 0.166 217 581k
MCMC (SfM) 28.79 0.888 0.190 157
EDGS 29.07 0.896 0.163 188
Ours ($\beta=0.04$) 29.32 0.898 0.157 116 927k
MCMC (SfM) 29.05 0.895 0.177 105
EDGS 29.35 0.902 0.151 134
Ours 29.60 0.904 0.144 114 1.75M
MCMC 29.46 0.904 0.157 89
EDGS 29.85 0.909 0.137 99
3DGS 29.30 0.896 0.171 88
Table A2: No-budget OMMO. Comparison of reconstruction quality without a fixed limit on the number of primitives. For each cell, our method is run with a different $\beta$ parameter, which determines the number of Gaussians generated, and the other methods are limited to this number. In the last cell, the number of Gaussians is set to match the amount produced by 3DGS, with all other methods constrained accordingly. Results are averaged over the OMMO [33] dataset.
Mip-NeRF360 [3] OMMO [33] Tanks & Temples [26] DeepBlending [19]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Number of Gaussians limited to 11M
3DGS [24] (SfM init.) 28.71 0.846 0.222 29.40 0.894 0.178 23.60 0.826 0.245 28.93 0.881 0.292
Foroutan et al. [14] 29.22 0.876 0.177 29.32 0.896 0.163 23.75 0.829 0.230 29.60 0.886 0.282
MCMC [25] (rand. init.) 28.98 0.865 0.204 28.83 0.888 0.185 23.41 0.824 0.249 28.94 0.877 0.303
MCMC [25] (SfM init.) 29.23 0.875 0.190 29.18 0.895 0.174 23.93 0.836 0.233 29.67 0.887 0.288
MCMC [25] (iNGP init.) 29.32 0.879 0.173 28.76 0.887 0.181 23.37 0.804 0.268 29.87 0.892 0.276
GaussianPro [9] 28.56 0.845 0.226 28.79 0.885 0.189 23.01 0.818 0.256 29.27 0.887 0.291
Perceptual-GS [64] 29.27 0.876 0.176 29.00 0.888 0.179 23.26 0.828 0.240 29.41 0.889 0.286
EDGS [27] 29.18 0.877 0.168 29.58 0.903 0.149 23.66 0.842 0.200 29.43 0.889 0.271
Ours 29.37 0.880 0.168 29.44 0.899 0.154 23.70 0.835 0.207 29.69 0.889 0.273
  • $\dagger$
    Original code was not publicly available. Our implementation uses iNGP initialization and does not include the additional depth-based loss.
Table A3: Quantitative results with a Gaussian number limit of 1M. We highlight the best, second best and third best results among methods with comparable numbers of Gaussians.
Refer to caption Refer to caption
Refer to caption Refer to caption
Ours Mini-Splatting2 [11]
Figure A1: Qualitative comparison between our approach and Mini-Splatting2 [11] on the garden scene from Mip-NeRF360 [3] and on the train scene from Tanks & Temples [26].
Table A1 and Table A2 present results for the no-budget scenario, which were used for plots in Figure 6. We use various values of the $\beta$ parameter without specifying a budget for our method, except for the last cell, which uses the budget set by the number of Gaussians generated by 3D Gaussian Splatting. For each cell, we also compare MCMC [25] and EDGS [27] where they are set to match the number of Gaussians produced in each of the cells. The comparisons show that our method is able to produce high-quality results even without specifying a budget, instead adjusting the number of primitives based on the scene complexity, while still remaining sparse in the number of primitives. Notably, even when setting $\beta=0$, which effectively disables densification, our method still performs well due to a dense initialization and effective filtering of unnecessary primitives enforced by the opacity penalty, consistent with findings in  [27]. However, this configuration does not allow explicit control over the number of primitives and may limit the achievable reconstruction quality.
If an even lower number of Gaussians is desired while still using a no-budget scenario, the number of initialized Gaussians can be reduced, effectively lowering the number of primitives generated at the end.
We show in Table A3 the results on the selected benchmarks for the budget of 1M primitives. Additionally, we present per-scene results for 100k Gaussians in Table A8, 500k Gaussians in Table A9, 1M Gaussians in Table A10, and 2M Gaussians in Table A11. We also show additional qualitative results with Mini-Splatting2 [11] in Figure A1.
Appendix A2 Additional ablations and comparisons
Ours 3DGS (SfM) MCMC (SfM) EDGS Perceptual-GS
# Memory (MiB) 9545 9049 8671 14517 12403
Table A4: Peak GPU memory usage on the 15 scene from the OMMO dataset [33] on the maximum budget of 500k primitives. We highlight the best, second best and third best results among all.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ Train (min) $\downarrow$
ConeGS 29.14 0.892 0.170 217 25
ConeGS (iNGP retrain)
29.27 0.893 0.168 212 29
Table A5: Ablation on additional iNGP training on the OMMO [33] dataset with 500k primitives, evaluating the effect of continuing training iNGP in parallel with the full 3DGS optimization. We highlight the best, second best and third best results among all.
[…]
Figure A2: Number of primitives pruned, added, accumulated, as well as the total number, during each densification, on different budget scenarios. Results obtained from the garden scene from Mip-NeRF360 [3].
Mip-NeRF360 [3] OMMO [33] Tanks & Temples [26] DeepBlending [19]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Number of Gaussians limited to 11M
Ours 29.37 0.880 0.168 29.44 0.899 0.154 23.70 0.835 0.207 29.69 0.889 0.273
Ours, iNGP training during 3DGS 29.25 0.879 0.168 29.54 0.900 0.152 23.72 0.836 0.207 29.73 0.889 0.273
Table A6: Ablation on additional iNGP training for the budget of 1M primitives. We highlight the best, second best and third best results among methods with comparable numbers of Gaussians.
We evaluate the peak GPU memory usage during optimization for several methods in Table A4. The results show that our method is very close to 3DGS [24] and MCMC [25] in terms of memory consumption, while remaining considerably lower than EDGS [27] and Perceptual-GS [64]. This efficiency allows our method to run on a wider range of GPUs, making it more accessible to devices with limited memory.
We expand on ablation (c) from the main paper, where iNGP continues training in parallel with the full 3DGS optimization. Since iNGP training, like densification, is guided by the L1 error from 3DGS, the iNGP model can better focus on regions where 3DGS reconstruction may fall short, potentially improving densification. The original ablation was tested on a budget of 100k primitives, which may not fully reveal this effect. To explore further, we run experiments with larger budgets. On the 500k budget for challenging OMMO [33] scenes (Table A5), we observe additional performance gains, though at the cost of longer training time. We also test this setup with a budget of 1M primitives across all datasets (Table A6), finding minimal improvements on difficult scenes and slight decreases on Mip-NeRF360, which may be caused by overfitting to certain areas.
Figure A2 illustrates the number of primitives added and removed during each densification and pruning step. It also tracks the accumulation buffer of primitives. When no budget is specified, all accumulated primitives are added to the scene (see Eq. 17). Under a fixed budget, however, not all of them are used. This is because accumulation happens every iteration and must remain available for pruning and densification, while the exact number that can be added is unknown beforehand. As a result, the system accumulates more primitives than are usually required to keep the total count close to the budget after pruning (see Eq. 16).
The primitive size is defined to be approximately one pixel from a single viewpoint. From other viewpoints, this size is not exactly one pixel, but should remain close. To confirm this, we measure the size of newly added pixel-sized primitives from multiple viewpoints and report the results in Figure A3. The distribution shows that the apparent size remains near one pixel, with very few cases above four pixels or below 0.2 pixels.
[…]
Figure A3: The histogram of perceived image-space size of pixel-sized Gaussians rendered from different viewpoints. Size expressed in pixel widths. Analysis doesn’t include the low-pass filter from 3DGS rasterization.
Appendix A3 Scene structure
[…]
Figure A4: Histograms of the Gaussian scaling values for our method and MCMC with SfM initialization on 2M Gaussians. The red lines indicate the minimum scaling required for a Gaussian to cover at least one pixel, disregarding the low-pass filter.
Ours
MCMC
Figure A5: Shrunk Gaussians. Visual comparison between the proposed method and MCMC, rendered using Gaussians scaled to half their original size. Both methods were trained with a limit of 1 million primitives on the bicycle and bonsai scenes from Mip-NeRF360 [3], and the 10 scene from OMMO [33].
Ours MCMC (SfM) MCMC 3DGS (SfM)
# Gaussians per pixel 30.72 49.55 48.45 34.74
Table A7: Blending analysis. Mean number of Gaussians contributing to alpha blending per pixel, averaged across the Mip-NeRF360 [3] dataset using a budget of 1M Gaussians. We highlight the best, second best and third best results among all.
To confirm that our method produces reconstructions with desirable characteristics, such as a balanced distribution of scaling values and placement close to surfaces, which are often important for downstream tasks, we analyze the scenes created with our method in comparison to MCMC [25].
Although Gaussian Splatting with the MCMC densification strategy explores the scene effectively, it also has clear shortcomings. Its cloning strategy and scaling penalty often cause Gaussians to shrink strongly along certain dimensions. The low-pass filter used by 3DGS renderer can hide this effect visually, but it prevents Gaussians from expanding in those directions and interferes with tasks that depend on accurate scales, such as MCMC position error. As shown in Figure A4, many of its Gaussians shrink below the pixel radius, in contrast to the more balanced distribution produced by our method.
Figure A5 shows that our approach produces Gaussians that are more uniform in size and placed closer to object surfaces. This avoids an overreliance on oversized background primitives located far from the surface. This primitive distribution not only improves geometric alignment but also increases rendering speed by reducing blending and sorting overhead during rasterization. To support this hypothesis, Table A7 reports the mean number of Gaussians blended per pixel. Our method consistently requires less blending compared to the selected benchmarks. The difference is especially large compared to MCMC, which blends over $60\%$ more Gaussians per pixel on average.
Appendix A4 Failure cases
While our method generally produces strong reconstructions on almost all tested scenes, its reliance on iNGP can make it susceptible to floaters in challenging scenarios such as noisy poses, very large or sparse-view scenes, or other cases where reliable iNGP reconstruction is difficult. One such example is shown in Figure A6, where the scene contains distortions and lacks sufficient viewpoint coverage near the cameras. This leads to spurious high-density regions in iNGP close to the cameras, which in turn degrades densification quality by placing Gaussians in incorrect regions. Although this can reduce performance, the method still produces high-quality reconstructions overall (see per-scene results).
GT ConeGS iNGP
Figure A6: Depth maps. Depth maps for ConeGS and the iNGP reconstruction model on the 01 scene from the OMMO [33] dataset. Since the Gaussians are generated from an iNGP reconstruction, the presence of floaters in it leads to floaters also appearing in the Gaussian Splatting results.
Appendix A5 Initialization visualizations
The choice of initialization has a strong impact on the performance of 3DGS reconstruction. A sufficiently good initialization can even remove the need for additional densification ([27], Table A1, Table A2). We visualize different initialization types in Figure A7. Initialization from a sparse SfM point cloud often leaves large gaps that are difficult to fill correctly, while our initialization produces a more uniform coverage. Another important strategy is initialization based on pixel width. This approach does not provide a useful inductive bias of larger primitives and can lead to more gaps in unobserved regions, although it may offer faster rendering due to reduced blending (Section 5.2). Finally, our formulation allows scaling the initial primitives in image space, producing a balance between these two types of initialization.
SfM point cloud init.
Our init.
1× pixel size init.
10× pixel size init.
Figure A7: Initialization Methods Comparison. Visual comparison of four initialization methods for 3DGS reconstruction on the bicycle scene from Mip-NeRF 360 dataset [3]: (1) SfM point cloud initialization, (2) iNGP initialization with kNN-based sizing, (3) iNGP initialization with 1× sizing, and (4) iNGP initialization using the smaller of 10× pixel size or kNN-based sizing.
Appendix A6 Gaussian-Based Radiance Field training
To establish a more direct connection between the volumetric ray marching procedure in Neural Radiance Fields (NeRF) and the rendering in 3D Gaussian Splatting (3DGS), the set of conical frustums sampled along a cone in Mip-NeRF and its derivatives [2, 3, 4] can be reinterpreted as a set of 3D Gaussian primitives, covering approximately the same area. The purpose of this reformulation is to enable training a neural radiance field model using only the 3DGS renderer, without using the traditional volumetric ray integration.
Figure A8: Training equivalence. Illustration of the equivalence between training an implicit radiance field model using NeRF-style ray marching and 3D Gaussian Splatting rasterization.
To align the 3DGS and NeRF rendering formulations, it is necessary to demonstrate that an entire ray can be equivalently represented as a sequence of Gaussians such that, when rendered, the result is equal to the volumetric integration process. This is achieved by casting cones along pixel directions and subdividing each into a set of conical frustums. Each frustum is then mapped to a Gaussian primitive, where the midpoint
$t_{\mu,i}=\frac{t_{i}+t_{i+1}}{2}$ (A1)
is used to define the Gaussian center $\mathbf{p}_{i}$. The view-dependent RGB color $\mathbf{c}_{i}$ and scalar density $\sigma_{i}$, which is converted to opacity $o_{i}$ using Eq. 6, are predicted by a neural network. As detailed in Section 4.1, the predicted RGB color $\mathbf{c}_{i}$ is encoded as spherical harmonics coefficients, which are compatible with the 3DGS rasterizer.
Since the Gaussians lie along the cone and their centers are co-linear with the pixel center and the camera origin, their projected 2D means fall directly on the target pixel. Consequently, the maximum opacity contribution aligns with the pixel center, and the kernel evaluation satisfies
$K(\mathbf{p}_{c},\boldsymbol{\mu}^{\text{2D}}_{i},\boldsymbol{\Sigma}^{\text{2D}}_{i})=1.$ (A2)
This condition ensures that the Gaussian opacity $o_{i}$ corresponds directly to the opacity $\alpha_{i}$ used in blending operations, as defined in Eq. 6 and Eq. 2. The opacity is therefore determined by the predicted density and the length of the corresponding frustum.
To maintain consistent 2D projection footprints across different depths, each Gaussian’s 3D covariance must be adjusted based on its location along the cone. Because elongating Gaussians along the ray direction does not affect the resulting 2D projection, the scaling can be derived using the pixel footprint size and set isotropically using the formulation in Eq. 15, while the quaternion can be set to the identity quaternion.
Given these assignments for position, color, opacity, scale, and orientation, each Gaussian can be rendered using the 3DGS renderer to produce the same pixel color as that produced by volumetric rendering. Assuming no low-pass filtering is applied and that only a single image is rendered at a time, it becomes possible to train an Instant-NGP-style model using only the 3DGS rendering pipeline. Although this approach is marginally slower than traditional ray marching, it yields equivalent radiance field representations. A visual comparison of both approaches is shown in Figure A8.
This reinterpretation also provides a more principled foundation for the proposed densification strategy. Instead of using depth, the strategy can be understood as selecting a Gaussian that corresponds to a surface-level conical frustum of pixels with high photometric error. This establishes a clearer theoretical link between cone-based densification and neural implicit models such as NeRF.
Scene Ours 3DGS (SfM) MCMC (SfM) EDGS
Mip-NeRF360 [3]
Bicycle 24.05 / 0.652 / 0.379 / 368 20.99 / 0.470 / 0.543 / 466 23.34 / 0.630 / 0.400 / 339 23.68 / 0.635 / 0.393 / 444
Garden 24.69 / 0.709 / 0.336 / 431 22.87 / 0.589 / 0.445 / 430 24.51 / 0.710 / 0.350 / 569 24.47 / 0.706 / 0.341 / 493
Stump 25.71 / 0.707 / 0.351 / 456 22.79 / 0.547 / 0.498 / 506 25.16 / 0.680 / 0.370 / 402 25.08 / 0.673 / 0.377 / 588
Room 31.09 / 0.903 / 0.254 / 275 27.02 / 0.859 / 0.328 / 267 30.37 / 0.900 / 0.270 / 224 30.08 / 0.896 / 0.265 / 317
Counter 28.22 / 0.879 / 0.251 / 229 25.64 / 0.845 / 0.307 / 252 27.83 / 0.880 / 0.260 / 198 27.77 / 0.877 / 0.251 / 245
Kitchen 29.73 / 0.896 / 0.183 / 239 20.53 / 0.671 / 0.440 / 297 28.37 / 0.890 / 0.210 / 246 28.84 / 0.891 / 0.186 / 269
Bonsai 30.67 / 0.920 / 0.241 / 276 25.41 / 0.866 / 0.331 / 326 29.84 / 0.910 / 0.260 / 281 29.73 / 0.906 / 0.256 / 298
Average 27.74 / 0.809 / 0.285 / 325 23.61 / 0.692 / 0.413 / 363 27.06 / 0.800 / 0.303 / 323 27.09 / 0.798 / 0.296 / 379
OMMO [33]
01 22.58 / 0.625 / 0.458 / 212 20.53 / 0.543 / 0.575 / 159 22.66 / 0.620 / 0.470 / 180 22.50 / 0.608 / 0.475 / 222
03 25.39 / 0.842 / 0.244 / 356 24.09 / 0.803 / 0.295 / 112 24.30 / 0.810 / 0.290 / 217 24.85 / 0.827 / 0.259 / 313
05 27.95 / 0.863 / 0.246 / 415 26.87 / 0.841 / 0.296 / 354 27.70 / 0.860 / 0.270 / 318 27.69 / 0.856 / 0.259 / 398
06 27.30 / 0.913 / 0.202 / 321 26.53 / 0.899 / 0.229 / 107 26.25 / 0.910 / 0.210 / 339 26.60 / 0.910 / 0.198 / 339
10 29.31 / 0.853 / 0.252 / 338 28.86 / 0.833 / 0.291 / 125 28.70 / 0.840 / 0.280 / 337 28.78 / 0.839 / 0.275 / 341
13 30.36 / 0.899 / 0.211 / 449 29.79 / 0.889 / 0.247 / 171 29.45 / 0.880 / 0.250 / 378 28.98 / 0.878 / 0.234 / 468
14 29.73 / 0.924 / 0.145 / 394 27.09 / 0.874 / 0.220 / 237 29.26 / 0.920 / 0.160 / 360 28.92 / 0.909 / 0.169 / 374
15 28.10 / 0.893 / 0.182 / 405 27.80 / 0.877 / 0.213 / 153 27.72 / 0.890 / 0.200 / 427 27.58 / 0.877 / 0.215 / 420
Average 27.59 / 0.852 / 0.242 / 361 26.45 / 0.820 / 0.296 / 177 27.00 / 0.841 / 0.266 / 320 26.99 / 0.838 / 0.261 / 359
Tanks & Temples [26]
Truck 24.54 / 0.822 / 0.291 / 184 23.59 / 0.803 / 0.318 / 128 23.86 / 0.810 / 0.316 / 155 23.33 / 0.799 / 0.314 / 183
Train 21.70 / 0.759 / 0.328 / 204 21.17 / 0.744 / 0.349 / 118 21.13 / 0.749 / 0.347 / 160 21.31 / 0.756 / 0.334 / 180
Average 23.12 / 0.790 / 0.309 / 194 22.38 / 0.774 / 0.334 / 123 22.49 / 0.780 / 0.332 / 158 22.32 / 0.778 / 0.324 / 182
Deep Blending [19]
Dr Johnson 28.78 / 0.875 / 0.338 / 447 23.58 / 0.808 / 0.440 / 521 28.16 / 0.869 / 0.343 / 550 27.66 / 0.865 / 0.347 / 444
Playroom 30.09 / 0.885 / 0.318 / 489 25.72 / 0.845 / 0.384 / 489 29.72 / 0.883 / 0.323 / 552 29.19 / 0.879 / 0.327 / 486
Average 29.44 / 0.880 / 0.328 / 468 24.65 / 0.827 / 0.412 / 505 28.94 / 0.876 / 0.333 / 551 28.43 / 0.872 / 0.337 / 465
Table A8: Detailed results on a selection of datasets and methods with the number of Gaussians limited to 100k. Each field contains PSNR, SSIM, LPIPS and FPS respectively. We highlight the best, second best and third best results among all. Slight discrepancies from the main table are due to rounding.
Scene Ours 3DGS (SfM) MCMC (SfM) EDGS
Mip-NeRF360 [3]
Bicycle 25.25 / 0.758 / 0.242 / 234 24.06 / 0.651 / 0.370 / 174 24.90 / 0.740 / 0.282 / 162 24.95 / 0.751 / 0.251 / 248
Garden 26.87 / 0.837 / 0.157 / 246 25.16 / 0.743 / 0.296 / 297 26.36 / 0.823 / 0.189 / 221 26.49 / 0.833 / 0.161 / 252
Stump 27.03 / 0.793 / 0.215 / 255 25.27 / 0.688 / 0.348 / 226 26.81 / 0.779 / 0.250 / 175 26.50 / 0.768 / 0.237 / 296
Room 32.02 / 0.924 / 0.205 / 192 31.47 / 0.913 / 0.234 / 112 31.72 / 0.921 / 0.221 / 114 31.36 / 0.923 / 0.203 / 168
Counter 28.87 / 0.906 / 0.194 / 148 28.86 / 0.901 / 0.213 / 113 28.95 / 0.907 / 0.205 / 93 29.05 / 0.912 / 0.184 / 137
Kitchen 31.38 / 0.929 / 0.124 / 172 30.76 / 0.918 / 0.143 / 111 31.04 / 0.921 / 0.141 / 111 31.35 / 0.928 / 0.122 / 154
Bonsai 32.16 / 0.944 / 0.191 / 190 31.95 / 0.936 / 0.218 / 135 31.97 / 0.939 / 0.210 / 128 32.02 / 0.942 / 0.191 / 170
Average 29.08 / 0.870 / 0.190 / 205 28.22 / 0.821 / 0.260 / 167 28.82 / 0.861 / 0.214 / 143 28.82 / 0.865 / 0.193 / 204
OMMO [33]
01 23.46 / 0.690 / 0.356 / 140 23.81 / 0.682 / 0.385 / 102 23.73 / 0.685 / 0.378 / 90 23.63 / 0.686 / 0.363 / 84
03 27.01 / 0.887 / 0.181 / 197 26.18 / 0.868 / 0.220 / 111 26.50 / 0.875 / 0.212 / 105 27.16 / 0.892 / 0.176 / 93
05 28.61 / 0.877 / 0.199 / 232 28.53 / 0.874 / 0.235 / 162 28.71 / 0.877 / 0.233 / 154 28.79 / 0.881 / 0.196 / 122
06 27.79 / 0.932 / 0.159 / 195 27.59 / 0.931 / 0.161 / 120 26.99 / 0.934 / 0.159 / 152 27.38 / 0.941 / 0.135 / 105
10 31.16 / 0.905 / 0.165 / 235 31.04 / 0.894 / 0.194 / 143 30.56 / 0.892 / 0.192 / 168 31.11 / 0.906 / 0.165 / 129
13 33.02 / 0.949 / 0.115 / 280 32.40 / 0.939 / 0.149 / 165 32.18 / 0.934 / 0.155 / 175 31.95 / 0.944 / 0.123 / 151
14 31.57 / 0.950 / 0.096 / 225 31.57 / 0.946 / 0.107 / 135 31.14 / 0.946 / 0.108 / 130 31.51 / 0.950 / 0.095 / 112
15 30.50 / 0.944 / 0.090 / 231 30.59 / 0.934 / 0.114 / 149 29.92 / 0.933 / 0.113 / 162 30.52 / 0.942 / 0.095 / 125
Average 29.14 / 0.892 / 0.170 / 217 28.96 / 0.884 / 0.196 / 136 28.72 / 0.885 / 0.194 / 142 29.01 / 0.893 / 0.169 / 115
Tanks & Temples [26]
Truck 25.41 / 0.858 / 0.204 / 122 24.80 / 0.842 / 0.253 / 93 25.24 / 0.851 / 0.242 / 68 24.66 / 0.850 / 0.212 / 114
Train 21.97 / 0.799 / 0.253 / 140 22.28 / 0.790 / 0.277 / 72 22.12 / 0.798 / 0.275 / 86 22.27 / 0.812 / 0.246 / 104
Average 23.69 / 0.829 / 0.229 / 131 23.54 / 0.816 / 0.265 / 82 23.68 / 0.825 / 0.259 / 77 23.47 / 0.831 / 0.229 / 109
Deep Blending [19]
Dr Johnson 29.20 / 0.890 / 0.291 / 270 28.85 / 0.881 / 0.307 / 151 28.71 / 0.884 / 0.313 / 180 28.60 / 0.888 / 0.292 / 269
Playroom 30.51 / 0.892 / 0.279 / 287 29.66 / 0.883 / 0.298 / 161 30.16 / 0.889 / 0.303 / 209 30.06 / 0.890 / 0.281 / 277
Average 29.86 / 0.891 / 0.285 / 278 29.26 / 0.882 / 0.303 / 156 29.44 / 0.887 / 0.308 / 194 29.33 / 0.889 / 0.287 / 273
Table A9: Detailed results on a selection of datasets and methods with the number of Gaussians limited to 500k. Each field contains PSNR, SSIM, LPIPS and FPS respectively. We highlight the best, second best and third best results among all. Slight discrepancies from the main table are due to rounding.
Scene Ours 3DGS (SfM) MCMC (SfM) EDGS
Mip-NeRF360 [3]
Bicycle 25.45 / 0.778 / 0.198 / 156 24.40 / 0.688 / 0.325 / 129 25.34 / 0.768 / 0.238 / 106 25.26 / 0.778 / 0.199 / 161
Garden 27.42 / 0.861 / 0.115 / 162 26.70 / 0.827 / 0.172 / 142 26.83 / 0.847 / 0.147 / 135 27.05 / 0.857 / 0.118 / 158
Stump 27.23 / 0.802 / 0.184 / 164 25.64 / 0.719 / 0.301 / 162 27.21 / 0.798 / 0.213 / 110 26.71 / 0.783 / 0.203 / 179
Room 32.30 / 0.928 / 0.196 / 139 31.62 / 0.918 / 0.221 / 82 32.09 / 0.926 / 0.208 / 79 31.71 / 0.928 / 0.191 / 111
Counter 29.21 / 0.911 / 0.181 / 107 29.11 / 0.907 / 0.201 / 75 29.24 / 0.913 / 0.191 / 61 29.27 / 0.917 / 0.171 / 94
Kitchen 31.66 / 0.932 / 0.117 / 118 31.18 / 0.924 / 0.131 / 78 31.44 / 0.927 / 0.130 / 72 31.83 / 0.933 / 0.114 / 101
Bonsai 32.33 / 0.945 / 0.183 / 132 32.34 / 0.941 / 0.205 / 95 32.46 / 0.944 / 0.199 / 81 32.44 / 0.947 / 0.181 / 115
Average 29.37 / 0.880 / 0.168 / 140 28.71 / 0.846 / 0.222 / 109 29.23 / 0.875 / 0.189 / 92 29.18 / 0.878 / 0.168 / 131
OMMO [33]
01 23.73 / 0.711 / 0.315 / 104 24.18 / 0.703 / 0.350 / 73 24.18 / 0.709 / 0.341 / 63 24.02 / 0.709 / 0.321 / 84
03 27.39 / 0.896 / 0.167 / 127 26.96 / 0.886 / 0.197 / 64 27.20 / 0.890 / 0.189 / 67 27.72 / 0.905 / 0.157 / 85
05 28.56 / 0.877 / 0.185 / 157 28.72 / 0.877 / 0.227 / 115 28.98 / 0.883 / 0.217 / 98 29.15 / 0.885 / 0.181 / 101
06 27.87 / 0.936 / 0.150 / 136 27.68 / 0.933 / 0.158 / 110 27.46 / 0.939 / 0.146 / 103 27.75 / 0.945 / 0.127 / 102
10 31.60 / 0.914 / 0.147 / 170 31.47 / 0.906 / 0.171 / 104 30.89 / 0.904 / 0.169 / 120 31.82 / 0.920 / 0.138 / 116
13 33.55 / 0.957 / 0.098 / 190 33.13 / 0.949 / 0.129 / 116 32.83 / 0.945 / 0.131 / 124 32.83 / 0.955 / 0.100 / 125
14 31.80 / 0.953 / 0.090 / 146 31.93 / 0.951 / 0.097 / 89 31.44 / 0.950 / 0.099 / 96 32.08 / 0.955 / 0.086 / 92
15 30.99 / 0.950 / 0.081 / 151 31.13 / 0.943 / 0.096 / 100 30.42 / 0.942 / 0.097 / 104 31.25 / 0.950 / 0.079 / 104
Average 29.44 / 0.899 / 0.154 / 148 29.40 / 0.894 / 0.178 / 96 29.18 / 0.895 / 0.174 / 97 29.58 / 0.903 / 0.149 / 101
Tanks & Temples [26]
Truck 25.34 / 0.864 / 0.181 / 97 25.06 / 0.851 / 0.234 / 70 25.57 / 0.862 / 0.217 / 45 25.02 / 0.862 / 0.181 / 82
Train 22.05 / 0.805 / 0.232 / 109 22.13 / 0.801 / 0.255 / 51 22.29 / 0.811 / 0.249 / 47 22.29 / 0.823 / 0.220 / 78
Average 23.70 / 0.835 / 0.207 / 103 23.60 / 0.826 / 0.245 / 60 23.93 / 0.837 / 0.233 / 46 23.66 / 0.843 / 0.201 / 80
Deep Blending [19]
Dr Johnson 28.98 / 0.886 / 0.282 / 194 28.85 / 0.884 / 0.293 / 106 29.06 / 0.884 / 0.291 / 164 28.75 / 0.890 / 0.277 / 180
Playroom 30.39 / 0.891 / 0.264 / 200 29.02 / 0.877 / 0.290 / 112 30.28 / 0.890 / 0.284 / 173 30.11 / 0.889 / 0.265 / 182
Average 29.69 / 0.889 / 0.273 / 197 28.94 / 0.881 / 0.292 / 109 29.67 / 0.887 / 0.288 / 168 29.43 / 0.890 / 0.271 / 181
Table A10: Detailed results on a selection of datasets and methods with the number of Gaussians limited to 1M. Each field contains PSNR, SSIM, LPIPS and FPS respectively. We highlight the best, second best and third best results among all. Slight discrepancies from the main table are due to rounding.
Scene Ours 3DGS (SfM) MCMC (SfM) EDGS
Mip-NeRF360 [3]
Bicycle 25.43 / 0.781 / 0.175 / 99 24.85 / 0.729 / 0.267 / 79 25.56 / 0.786 / 0.205 / 71 25.48 / 0.792 / 0.168 / 93
Garden 27.59 / 0.870 / 0.097 / 97 27.22 / 0.852 / 0.129 / 84 27.33 / 0.862 / 0.122 / 82 27.46 / 0.870 / 0.097 / 89
Stump 27.03 / 0.796 / 0.177 / 100 26.18 / 0.749 / 0.255 / 100 27.39 / 0.808 / 0.189 / 69 26.79 / 0.788 / 0.186 / 100
Room 32.32 / 0.929 / 0.189 / 92 31.80 / 0.919 / 0.218 / 63 32.14 / 0.929 / 0.199 / 53 31.89 / 0.930 / 0.184 / 81
Counter 29.23 / 0.912 / 0.175 / 70 29.07 / 0.907 / 0.201 / 68 29.41 / 0.917 / 0.181 / 42 29.37 / 0.918 / 0.164 / 65
Kitchen 31.80 / 0.934 / 0.114 / 72 31.60 / 0.927 / 0.126 / 54 32.00 / 0.931 / 0.122 / 47 32.06 / 0.935 / 0.111 / 65
Bonsai 32.63 / 0.948 / 0.177 / 84 32.34 / 0.941 / 0.204 / 84 32.80 / 0.948 / 0.189 / 53 32.62 / 0.947 / 0.175 / 85
Average 29.43 / 0.881 / 0.158 / 88 29.01 / 0.861 / 0.200 / 76 29.52 / 0.883 / 0.172 / 60 29.38 / 0.883 / 0.155 / 83
OMMO [33]
01 23.92 / 0.725 / 0.283 / 73 24.51 / 0.721 / 0.321 / 49 24.51 / 0.731 / 0.303 / 45 24.24 / 0.725 / 0.285 / 64
03 27.73 / 0.902 / 0.155 / 70 27.05 / 0.888 / 0.193 / 47 27.73 / 0.900 / 0.171 / 45 27.77 / 0.909 / 0.149 / 82
05 28.65 / 0.877 / 0.176 / 91 28.69 / 0.877 / 0.227 / 112 29.20 / 0.887 / 0.201 / 67 29.31 / 0.887 / 0.169 / 90
06 27.93 / 0.937 / 0.143 / 79 27.67 / 0.933 / 0.157 / 109 27.58 / 0.942 / 0.138 / 63 27.93 / 0.947 / 0.123 / 97
10 31.81 / 0.920 / 0.135 / 104 31.77 / 0.915 / 0.153 / 70 31.59 / 0.913 / 0.151 / 80 32.26 / 0.928 / 0.123 / 89
13 33.88 / 0.961 / 0.088 / 112 33.79 / 0.957 / 0.112 / 73 33.30 / 0.953 / 0.111 / 80 33.55 / 0.961 / 0.087 / 90
14 32.02 / 0.955 / 0.086 / 83 31.95 / 0.951 / 0.096 / 82 31.75 / 0.953 / 0.093 / 62 31.76 / 0.956 / 0.084 / 86
15 31.25 / 0.953 / 0.075 / 86 31.35 / 0.947 / 0.090 / 83 30.80 / 0.947 / 0.086 / 67 31.63 / 0.954 / 0.072 / 87
Average 29.65 / 0.904 / 0.143 / 87 29.60 / 0.899 / 0.169 / 78 29.56 / 0.903 / 0.157 / 64 29.81 / 0.908 / 0.137 / 86
Tanks & Temples [26]
Truck 25.43 / 0.865 / 0.167 / 69 25.39 / 0.859 / 0.217 / 53 25.86 / 0.869 / 0.193 / 38 25.09 / 0.865 / 0.162 / 55
Train 21.95 / 0.809 / 0.217 / 77 22.29 / 0.802 / 0.253 / 55 22.35 / 0.822 / 0.228 / 40 21.84 / 0.827 / 0.204 / 58
Average 23.69 / 0.837 / 0.192 / 73 23.84 / 0.831 / 0.235 / 54 24.11 / 0.846 / 0.211 / 39 23.47 / 0.846 / 0.183 / 56
Deep Blending [19]
Dr Johnson 29.01 / 0.884 / 0.274 / 127 28.75 / 0.886 / 0.282 / 77 28.83 / 0.882 / 0.285 / 104 28.65 / 0.887 / 0.268 / 113
Playroom 30.50 / 0.887 / 0.247 / 127 29.32 / 0.880 / 0.279 / 83 30.22 / 0.890 / 0.272 / 105 30.11 / 0.886 / 0.248 / 112
Average 29.76 / 0.886 / 0.261 / 127 29.04 / 0.883 / 0.281 / 80 29.53 / 0.886 / 0.279 / 104 29.38 / 0.887 / 0.258 / 112
Table A11: Detailed results on a selection of datasets and methods with the number of Gaussians limited to 2M. Each field contains PSNR, SSIM, LPIPS and FPS respectively. We highlight the best, second best and third best results among all. Slight discrepancies from the main table are due to rounding.